OpenStudy (anonymous):

Find the missing angle measure in triangleQRS. angleQ = ___°, angleR = 61°, angleS = 29°

OpenStudy (anonymous):

can i c a diagram

OpenStudy (anonymous):

thats all there was to the question

OpenStudy (anonymous):

Since the sum of angles are equal to 180 So We can write angleQ+angleR+angleS=180 so angQ+61+29=180 angQ+90=180 angQ=180-90 angQ=90 Answer 100 %

OpenStudy (anonymous):

i figured it out! thank you guys for trying

OpenStudy (anonymous):

Given: x + y = 15. Which of the following are ordered pairs of this equation? {(-3, 14),(0, 15),(3, 16)} {(-3, 16),(0, 15),(3, 14)} {(16, -3),(15, 0),(14, 3)}

OpenStudy (anonymous):

i need help with that roses